# Initialization using Update Approximation is a Silver Bullet for Extremely Efficient Low-Rank Fine-Tuning

### v3

@@ Appendix D Dataset Details
    HellaSwag (Zellers et al., 2019) challenges models to select the most plausible continuation of a given scenario from multiple possible endings.
@@
    ARC Easy (or ARC-e) (Clark et al., 2018) includes basic science questions at a grade-school level, offering simpler tasks to assess fundamental reasoning abilities.
@@
    PIQA (Bisk et al., 2020) evaluates physical commonsense reasoning, where models must choose the best action to take in a hypothetical scenario.
@@
    SIQA (Sap et al., 2019) tests social commonsense reasoning by asking models to predict the social consequences of human actions.
@@
    WinoGrande (Sakaguchi et al., 2021) presents sentence completion tasks requiring commonsense reasoning to select the correct binary option.
@@
    ARC Challenge (or ARC-c) (Clark et al., 2018) consists of more complex science questions designed to challenge models with sophisticated reasoning, beyond simple co-occurrence patterns.
@@
    OBQA (Mihaylov et al., 2018) features open-book, knowledge-intensive QA tasks that require multi-hop reasoning across multiple information sources.
@@
    BoolQ (Clark et al., 2019) involves answering yes/no questions based on real-world, naturally occurring queries.
The GLUE Benchmark is a comprehensive collection of tasks designed to evaluate natural language understanding (NLU) abilities. It included various datasets, including STS-B for measuring semantic textual similarity (Cer et al., 2017), RTE for recognizing textual entailment, MRPC for detecting paraphrases (Dolan & Brockett, 2005), CoLA for assessing linguistic acceptability (Warstadt et al., 2019), SST-2 for sentiment analysis (Socher et al., 2013), and QNLI for question-answer inference (Rajpurkar et al., 2018). GLUE’s broad scope makes it a standard benchmark for evaluating models like RoBERTa.

### v4

@@ Appendix I Dataset Details
    HellaSwag (51) challenges models to select the most plausible continuation of a given scenario from multiple possible endings.
@@
    ARC Easy (or ARC-e) (7) includes basic science questions at a grade-school level, offering simpler tasks to assess fundamental reasoning abilities.
@@
    PIQA (3) evaluates physical commonsense reasoning, where models must choose the best action to take in a hypothetical scenario.
@@
    SIQA (39) tests social commonsense reasoning by asking models to predict the social consequences of human actions.
@@
    WinoGrande (38) presents sentence completion tasks requiring commonsense reasoning to select the correct binary option.
@@
    ARC Challenge (or ARC-c) (7) consists of more complex science questions designed to challenge models with sophisticated reasoning, beyond simple co-occurrence patterns.
@@
    OBQA (31) features open-book, knowledge-intensive QA tasks that require multi-hop reasoning across multiple information sources.
@@
    BoolQ (6) involves answering yes/no questions based on real-world, naturally occurring queries.
The GLUE Benchmark is a comprehensive collection of tasks designed to evaluate natural language understanding (NLU) abilities. It included various datasets, including STS-B for measuring semantic textual similarity (5), RTE for recognizing textual entailment, MRPC for detecting paraphrases (11), CoLA for assessing linguistic acceptability (47), SST-2 for sentiment analysis (41), and QNLI for question-answer inference (36). GLUE’s broad scope makes it a standard benchmark for evaluating models like RoBERTa.
Appendix J Use of Large Language Models
LLMs are only used for small writing improvements, like polishing grammar and smoothing out phrasing.